# Numerical analysis for leaky-integrate-fire networks under Euler–Maruyama

### 11 Mar 2026 v1

@@ Abstract
For weak error we adapt a semigroup/backward-Kolmogorov argument to the reset map. The one-step defect splits into an interior Taylor term and a boundary-strip probability-flow term controlling missed and extra spikes, which yields global order $1$, $O(Th)$. We also derive a Lyapunov exponent formula coupling the stationary threshold flux to the reset saltation factor, and we outline deterministic and noisy recurrent extensions, including loop-truncated strong bounds governed by synaptic cycles and rate/density controls. The resulting trajectory-level theory separates single-trial spike-time fidelity from observable-level accuracy, making it relevant both for mechanistic neuroscience models and for neuromorphic AI systems that compute through sparse events and filtered readouts.
@@ 1.1 Motivation: why spiking resets change numerical analysis
We study current-based LIF networks with temporally exponential synapic kernels, discretized by Euler–Maruyama (EM). Because the voltage has no direct diffusion, a spike is a boundary hit for an advected voltage with random crossing speed $A=I-I_{\mathrm{th}}$. In fluctuation-driven regimes $A$ can be arbitrarily small. This forces a spike-time analysis, replaces the usual globally Lipschitz perturbation picture by a transversality problem at threshold, and produces logarithmic losses in strong error bounds. We start from a layered feedforward architecture, then extend our results to recurrent spiking networks.
@@ 1.2 Related work
At the single-neuron level, the spike time is a first-passage time. For diffusion-driven exits, Euler monitoring on a grid typically yields a $1/4$ strong exponent and requires boundary corrections for weak order one [13, 2, 1]. Our setting is degenerate: diffusion enters through $I$ but not through $v$. The singular factor is the inverse crossing speed at threshold, whose truncated moments diverge only logarithmically under a boundary-flux bound. This leads to mean-square strong error of size $h$ up to polylogarithmic factors, except for a terminal-time mismatch layer for discontinuous observables.
@@ 1.3 Contributions and interpretation
    When depth does not change the strong $h$-exponent. In noisy layers, the good-set recursion picks up one logarithmic inverse-moment factor per layer. If no intrinsic current noise enters layers $\ell\geq 2$ and downstream spikes have a deterministic transversality margin, then no new small-speed events are created downstream. The strong $h$-exponent is then inherited from layer $1$, and depth affects only the constants through weights and spike counts. Thus synaptic depth can amplify error constants without necessarily changing the basic step-size exponent.
@@
    Weak error is order $1$ with explicit constants. We obtain global weak order $1$ for spike-map compatible $\Phi\in C_{b}^{4}$ by a backward-Kolmogorov/Markov-operator argument that explicitly accounts for off-grid spikes. The one-step weak defect is $O(h^{2})$ and decomposes into diffusion truncation, synaptic timing/decay, and missed/extra spike decisions. The spike-decision term is controlled by a boundary-strip probability-flow estimate; no “spike snapping” is assumed. The global error is $O(Th)$ with constants explicit in firing-rate bounds, synaptic time scales, and weight norms; under uniform rate/weight bounds the depth dependence is at most linear. This is the regime most directly relevant to smooth population-level readouts.
@@ 1.4 Notation
$\Delta N^{h}_{\ell,j}(m)$ Numerical spikes of neuron $(\ell,j)$ in $t_{m},t_{m+1})$ (typically 0 or $1$).
@@
$S(I)$ Saltation (reset) factor at a spike with current $I$ (hybrid variational jump).
@@ 1.5 Organization
Section 2 introduces the model and EM discretization and summarizes the main results. Section 3 contains the STE-based strong analysis, including depth/time scaling and two regimes avoiding depth deterioration. Section 4 provides the weak order-$1$ proof with explicit constants. Section 5 derives Lyapunov exponents and their feedforward propagation, and discusses implications for long-time strong error. Sections 67 extend strong/weak bounds to recurrent networks under deterministic transversality and/or rate/density controls. The manuscript ends with a brief conclusion.
@@ 2.5 Main results
For noisy recurrent networks, under rate moment bounds ($R_{\mathrm{net}},R_{\mathrm{net},2}$) and boundary density controls ($\rho_{\max},\rho^{V}_{\max}$), we obtain pruning-type strong error bounds; on the strong good set, STE propagation admits an explicit loop-length truncation governed by the shortest directed cycle length of the synaptic graph (Theorem 6.13). Weak error remains order $1$ with constants depending explicitly on $R_{\mathrm{net}},R_{\mathrm{net},2}$ and $\|W\|_{1}$ (Theorem 7.2).
@@ 3.1 Warm-up: strong error when noise is injected directly into the voltage
For the current-based model studied below, $v$ has no diffusion. Therefore the “Brownian recrossing” mechanism behind the $O(Th^{1/2})$ mismatch probability in the warm-up disappears. Instead, the dominant strong obstruction becomes near-tangential crossings (small $A$), whose probability is governed by the boundary flux (13) and scales like $O(T\alpha^{2})$. False-positive numerical spikes remain possible (numerical fires but exact does not), but we show their probability is still only $O(Th^{1/2})$ under a mild voltage-density bound and is subleading in the depth recursion for $L\geq 2$ (Lemma 3.16 and Remark 3.18).
@@
3.2 Single-neuron strong STE bound (noise enters $I$ only)
@@
This is standard EM strong order $1/2$ for the OU current SDE and linear stability for the voltage ODE driven by $I$; see, e.g., [19]. ∎
@@ 3.7.4 Iterating the STE recursion
with the usual convention that an empty product equals $1$. Under the simplifying uniform assumptions
@@ 3.8.1 No intrinsic noise downstream
Theorem 3.31 (No downstream noise: the strong $h$-exponent is inherited from layer $1$).
@@
    Tangency-bad events can occur only in layer $1$ (downstream crossings have a deterministic margin):
@@
    Consequently, for each fixed depth $L$, the good-set spike-time error exponent is the one from layer $1$: if layer $1$ has current noise, then $\Delta_{L}(T)=O(\sqrt{h\,\log(1/h)})$ up to multiplicative constants; if layer $1$ is deterministic and mean-driven, then $\Delta_{L}(T)=O(h)$. The pruning bound for terminal-time observables inherits the same $h$-exponents, up to the mismatch probability from Lemma 3.16. Depth affects only multiplicative constants through weights and spike counts.
@@
If all layers are deterministic and have a uniform transversality margin, then one can take all $\alpha_{\ell}$ fixed and recover the deterministic (order-$1$) spike-time accuracy of forward Euler. In contrast, with Brownian current noise in a layer the crossing speed can be arbitrarily small with positive probability, so a.s. uniform transversality is generally unavailable; our $\alpha$-splitting framework is designed for this stochastic setting.
@@
4 Weak error: order $1$ with explicit dependence on depth, time, and weights
@@ 4.4 Quantities controlling the weak constant
Assume $u\in C^{1,4}([0,T)\times\mathsf{D})$ with bounded derivatives up to order $4$. Define
@@ 4.6 One-step and global weak error bounds
Theorem 4.4 (Global weak order $1$ with explicit TTLL scaling).
@@ 5.2.5 Spike-count bounds from input integrals
Between spikes, $x(t)$ increases at speed $\dot{x}=I(t)+f(x(t))\leq I(t)+f(0)$ since $f$ is nonincreasing. Each spike requires an increment of $1$ in $x$. Summing increments over interspike intervals yields the integral bound (74). ∎
@@ 5.3.2 Invariant measure and boundary flux
Assume that the joint process $(x(t),I(t))$ on $\{x<1\}$ with reinjection at $x=0$ is positive recurrent and admits a stationary ergodic distribution $\pi_{\infty}$ on $\mathsf{D}:=[0,1)\times\mathbb{R}$. Assume moreover that the stationary law admits a density $\rho_{\infty}(x,I)$ that is bounded in a neighborhood of $(1,I)$ for $I$ near the tangency set $\{f(1)+I=0\}$. In the Markov case (e.g. $I$ an OU diffusion), the corresponding stationary Fokker–Planck equation implies a stationary outflow boundary flux through $x=1$ given by
@@
More generally (beyond Markov), $J_{\infty}$ can be understood as the stationary Palm measure of $I$ at spike times; for our purposes the Markov/flux setting is sufficient.
@@ 5.5.1 A forcing estimate: spike-time divergence $\Rightarrow$ current divergence
Consider one presynaptic neuron (layer $1$) and one postsynaptic neuron (layer $2$) coupled by a weight $w$ with exponential synapse kernel $K(t;s)=e^{-(t-s)/\tau_{\mathrm{c}}}\mathbbm{1}_{\{t\geq s\}}$. Let $\{s_{1}^{k}\}$ and $\{\widetilde{s}_{1}^{k}\}$ be the spike times of the two copies of neuron $1$ (with matched indices), and define $\varepsilon_{1}^{k}:=\widetilde{s}_{1}^{k}-s_{1}^{k}$.
@@
Assume spike indices of neuron $1$ remain matched up to time $T$. Let $I_{2}^{\mathrm{ff}}$ and $\widetilde{I}_{2}^{\mathrm{ff}}$ be the feedforward parts of the postsynaptic currents driven by $N_{1}$ and $\widetilde{N}_{1}$. Then
@@ 5.5.2 A max-rule for layerwise Lyapunov exponents
For 1D IF driven by a common input and with $f^{\prime}\leq 0$, Lemma 5.5 implies that spike-count mismatch is uniformly bounded by $1$. This supports the matched-indices assumption above for single neurons and in regimes where effective order-preservation holds.
@@ 6.2 Model: recurrent current-based LIF network
The numerical method is Euler–Maruyama for $I$ and forward Euler for $v$, with grid-based spike detection and reset (as in Section 2).
@@ 6.3.2 Saltation matrix at a spike event
where $I_{p}^{-}=I_{p}(s^{-})$ and $e_{p}$ is the unit vector in the $v$-block at index $p$. Equivalently, acting on perturbations $(\delta v^{-},\delta I^{-})$,
@@ 6.3.3 Hybrid fundamental matrix and a Lyapunov upper bound
The saltation structure (93) shows that each spike of neuron $p$ injects its voltage perturbation $\delta v_{p}^{-}$ into downstream currents and voltages through the synaptic column $W_{:p}$ with amplification proportional to $1/A_{p}$. This is the mechanism underlying the cycle-gain criterion: if a directed cycle $C$ is repeatedly realized by the dynamics and the product of per-edge gains exceeds $1$, then the spike-time Lyapunov exponent is forced positive.
@@ 6.4.3 Tangency probability and spike-count mismatch probability
which treats false positives and false negatives symmetrically and ignores transient stepwise discrepancies that catch up before the observation time. Assume the local rate and burst controls (27)–(28) and the sub-Gaussian spike-time tail (29) hold for neuron $i$. Then Lemma 3.16 yields, for any $\delta\in(0,1]$,
@@ 6.4.4 Good-set STE inequality and explicit loop-length truncation
Fix truncation thresholds $\alpha_{i}\in(0,\alpha_{0}]$. Let $G(h,T;\alpha)$ be the event that all spikes up to $T$ satisfy $A_{i}^{k}\geq\alpha_{i}$ and that spike counts match at the observation time (no horizon mismatch), i.e. $N_{i}^{h}(T)=N_{i}(T)$ for every neuron $i$.
@@
Under Assumption 6.11, the remainder term $G^{K+1}\Delta$ is absent on the good set once $K\geq K_{\mathrm{eff}}(T)$: each multiplication by $G$ corresponds to taking one synaptic predecessor, so $(G^{K+1}\Delta)_{i}$ expands into contributions along directed, time-ordered influence chains of length $K+1$ ending at neuron $i$. By definition of $K_{\mathrm{eff}}(T)$, such chains do not exist on $G(h,T;\alpha)$ for $K=K_{\mathrm{eff}}(T)$. Therefore, on $G$ the Neumann series truncates and (104) holds. Taking $\|\cdot\|_{\infty}$ gives (105).
@@ 7.3 One-step weak defect: explicit dependence on $R_{\mathrm{net}}$ and $R_{\mathrm{net},2}$
7.4 Global weak order $1$ with explicit rate dependence
Theorem 7.2 (Recurrent global weak order $1$ with explicit $R_{\mathrm{net}},R_{\mathrm{net},2}$ scaling).
@@ 7.4 Global weak order $1$ with explicit rate dependence
The bursty synchronous regime corresponds to large $R_{\mathrm{net}}(T)$ and $R_{\mathrm{net},2}(T)$. The weak order remains $1$ under these bounds, but the constants can become large. Moreover, synchronized regimes can increase the strip-density constants $\rho^{V}_{\max,i}(T)$ and the velocity moments $A_{2,i}(T)$ (in small-noise limits these may blow up), further increasing $C_{\mathrm{weak}}^{\mathrm{rec}}(T,W)$.

### 12 Mar 2026 v2

@@ Abstract
For weak error we adapt a semigroup/backward-Kolmogorov argument to the reset map. The one-step defect splits into an interior Taylor term and a boundary-strip probability-flow term controlling missed and extra spikes, which yields global order 11, $O(Th)$. We also derive a Lyapunov exponent formula coupling the stationary threshold flux to the reset saltation factor, and we outline deterministic and noisy recurrent extensions, including loop-truncated strong bounds governed by synaptic cycles and rate/density controls. The resulting trajectory-level theory separates single-trial spike-time fidelity from observable-level accuracy, making it relevant both for mechanistic neuroscience models and for neuromorphic AI systems that compute through sparse events and filtered readouts.
@@ 1.1 Motivation: why spiking resets change numerical analysis
We study current-based LIF networks with temporally exponential synapic kernels, discretized by Euler–Maruyama (EM). Because the voltage has no direct diffusion, a spike is a boundary hit for an advected voltage with random crossing speed $A=I-I_{\mathrm{th}}$. In fluctuation-driven regimes AA can be arbitrarily small. This forces a spike-time analysis, replaces the usual globally Lipschitz perturbation picture by a transversality problem at threshold, and produces logarithmic losses in strong error bounds. We start from a layered feedforward architecture, then extend our results to recurrent spiking networks.
@@ 1.2 Related work
At the single-neuron level, the spike time is a first-passage time. For diffusion-driven exits, Euler monitoring on a grid typically yields a $1/4$ strong exponent and requires boundary corrections for weak order one [13, 2, 1]. Our setting is degenerate: diffusion enters through II but not through $v$. The singular factor is the inverse crossing speed at threshold, whose truncated moments diverge only logarithmically under a boundary-flux bound. This leads to mean-square strong error of size $h$ up to polylogarithmic factors, except for a terminal-time mismatch layer for discontinuous observables.
@@ 1.3 Contributions and interpretation
    When depth does not change the strong $h$-exponent. In noisy layers, the good-set recursion picks up one logarithmic inverse-moment factor per layer. If no intrinsic current noise enters layers $\ell\geq 2$ and downstream spikes have a deterministic transversality margin, then no new small-speed events are created downstream. The strong $h$-exponent is then inherited from layer 11, and depth affects only the constants through weights and spike counts. Thus synaptic depth can amplify error constants without necessarily changing the basic step-size exponent.
@@
    Weak error is order 11 with explicit constants. We obtain global weak order 11 for spike-map compatible $\Phi\in C_{b}^{4}$ by a backward-Kolmogorov/Markov-operator argument that explicitly accounts for off-grid spikes. The one-step weak defect is $O(h^{2})$ and decomposes into diffusion truncation, synaptic timing/decay, and missed/extra spike decisions. The spike-decision term is controlled by a boundary-strip probability-flow estimate; no “spike snapping” is assumed. The global error is $O(Th)$ with constants explicit in firing-rate bounds, synaptic time scales, and weight norms; under uniform rate/weight bounds the depth dependence is at most linear. This is the regime most directly relevant to smooth population-level readouts.
@@ 1.4 Notation
$\Delta N^{h}_{\ell,j}(m)$ Numerical spikes of neuron $(\ell,j)$ in $t_{m},t_{m+1})$ (typically 0 or 11).
@@
$S(I)$ Saltation (reset) factor at a spike with current II (hybrid variational jump).
@@ 1.5 Organization
Section 2 introduces the model and EM discretization and summarizes the main results. Section 3 contains the STE-based strong analysis, including depth/time scaling and two regimes avoiding depth deterioration. Section 4 provides the weak order-11 proof with explicit constants. Section 5 derives Lyapunov exponents and their feedforward propagation, and discusses implications for long-time strong error. Sections 67 extend strong/weak bounds to recurrent networks under deterministic transversality and/or rate/density controls. The manuscript ends with a brief conclusion.
@@ 2.5 Main results
For noisy recurrent networks, under rate moment bounds ($R_{\mathrm{net}},R_{\mathrm{net},2}$) and boundary density controls ($\rho_{\max},\rho^{V}_{\max}$), we obtain pruning-type strong error bounds; on the strong good set, STE propagation admits an explicit loop-length truncation governed by the shortest directed cycle length of the synaptic graph (Theorem 6.13). Weak error remains order 11 with constants depending explicitly on $R_{\mathrm{net}},R_{\mathrm{net},2}$ and $\|W\|_{1}$ (Theorem 7.2).
@@ 3.1 Warm-up: strong error when noise is injected directly into the voltage
For the current-based model studied below, $v$ has no diffusion. Therefore the “Brownian recrossing” mechanism behind the $O(Th^{1/2})$ mismatch probability in the warm-up disappears. Instead, the dominant strong obstruction becomes near-tangential crossings (small AA), whose probability is governed by the boundary flux (13) and scales like $O(T\alpha^{2})$. False-positive numerical spikes remain possible (numerical fires but exact does not), but we show their probability is still only $O(Th^{1/2})$ under a mild voltage-density bound and is subleading in the depth recursion for $L\geq 2$ (Lemma 3.16 and Remark 3.18).
@@
3.2 Single-neuron strong STE bound (noise enters II only)
@@
This is standard EM strong order $1/2$ for the OU current SDE and linear stability for the voltage ODE driven by II; see, e.g., [19]. ∎
@@ 3.7.4 Iterating the STE recursion
with the usual convention that an empty product equals 11. Under the simplifying uniform assumptions
@@ 3.8.1 No intrinsic noise downstream
Theorem 3.31 (No downstream noise: the strong $h$-exponent is inherited from layer 11).
@@
    Tangency-bad events can occur only in layer 11 (downstream crossings have a deterministic margin):
@@
    Consequently, for each fixed depth $L$, the good-set spike-time error exponent is the one from layer 11: if layer 11 has current noise, then $\Delta_{L}(T)=O(\sqrt{h\,\log(1/h)})$ up to multiplicative constants; if layer 11 is deterministic and mean-driven, then $\Delta_{L}(T)=O(h)$. The pruning bound for terminal-time observables inherits the same $h$-exponents, up to the mismatch probability from Lemma 3.16. Depth affects only multiplicative constants through weights and spike counts.
@@
If all layers are deterministic and have a uniform transversality margin, then one can take all $\alpha_{\ell}$ fixed and recover the deterministic (order-11) spike-time accuracy of forward Euler. In contrast, with Brownian current noise in a layer the crossing speed can be arbitrarily small with positive probability, so a.s. uniform transversality is generally unavailable; our $\alpha$-splitting framework is designed for this stochastic setting.
@@
4 Weak error: order 11 with explicit dependence on depth, time, and weights
@@ 4.4 Quantities controlling the weak constant
Assume $u\in C^{1,4}([0,T)\times\mathsf{D})$ with bounded derivatives up to order 44. Define
@@ 4.6 One-step and global weak error bounds
Theorem 4.4 (Global weak order 11 with explicit TTLL scaling).
@@ 5.2.5 Spike-count bounds from input integrals
Between spikes, $x(t)$ increases at speed $\dot{x}=I(t)+f(x(t))\leq I(t)+f(0)$ since $f$ is nonincreasing. Each spike requires an increment of 11 in $x$. Summing increments over interspike intervals yields the integral bound (74). ∎
@@ 5.3.2 Invariant measure and boundary flux
Assume that the joint process $(x(t),I(t))$ on $\{x<1\}$ with reinjection at $x=0$ is positive recurrent and admits a stationary ergodic distribution $\pi_{\infty}$ on $\mathsf{D}:=[0,1)\times\mathbb{R}$. Assume moreover that the stationary law admits a density $\rho_{\infty}(x,I)$ that is bounded in a neighborhood of $(1,I)$ for II near the tangency set $\{f(1)+I=0\}$. In the Markov case (e.g. II an OU diffusion), the corresponding stationary Fokker–Planck equation implies a stationary outflow boundary flux through $x=1$ given by
@@
More generally (beyond Markov), $J_{\infty}$ can be understood as the stationary Palm measure of II at spike times; for our purposes the Markov/flux setting is sufficient.
@@ 5.5.1 A forcing estimate: spike-time divergence $\Rightarrow$ current divergence
Consider one presynaptic neuron (layer 11) and one postsynaptic neuron (layer 22) coupled by a weight $w$ with exponential synapse kernel $K(t;s)=e^{-(t-s)/\tau_{\mathrm{c}}}\mathbbm{1}_{\{t\geq s\}}$. Let $\{s_{1}^{k}\}$ and $\{\widetilde{s}_{1}^{k}\}$ be the spike times of the two copies of neuron 11 (with matched indices), and define $\varepsilon_{1}^{k}:=\widetilde{s}_{1}^{k}-s_{1}^{k}$.
@@
Assume spike indices of neuron 11 remain matched up to time $T$. Let $I_{2}^{\mathrm{ff}}$ and $\widetilde{I}_{2}^{\mathrm{ff}}$ be the feedforward parts of the postsynaptic currents driven by $N_{1}$ and $\widetilde{N}_{1}$. Then
@@ 5.5.2 A max-rule for layerwise Lyapunov exponents
For 1D IF driven by a common input and with $f^{\prime}\leq 0$, Lemma 5.5 implies that spike-count mismatch is uniformly bounded by 11. This supports the matched-indices assumption above for single neurons and in regimes where effective order-preservation holds.
@@ 6.2 Model: recurrent current-based LIF network
The numerical method is Euler–Maruyama for II and forward Euler for $v$, with grid-based spike detection and reset (as in Section 2).
@@ 6.3.2 Saltation matrix at a spike event
where $I_{p}^{-}=I_{p}(s^{-})$ and $e_{p}$ is the unit vector in the $v$-block at index pp. Equivalently, acting on perturbations $(\delta v^{-},\delta I^{-})$,
@@ 6.3.3 Hybrid fundamental matrix and a Lyapunov upper bound
The saltation structure (93) shows that each spike of neuron $p$ injects its voltage perturbation $\delta v_{p}^{-}$ into downstream currents and voltages through the synaptic column $W_{:p}$ with amplification proportional to $1/A_{p}$. This is the mechanism underlying the cycle-gain criterion: if a directed cycle $C$ is repeatedly realized by the dynamics and the product of per-edge gains exceeds 11, then the spike-time Lyapunov exponent is forced positive.
@@ 6.4.3 Tangency probability and spike-count mismatch probability
which treats false positives and false negatives symmetrically and ignores transient stepwise discrepancies that catch up before the observation time. Assume the local rate and burst controls (27)–(28) and the sub-Gaussian spike-time tail (29) hold for neuron ii. Then Lemma 3.16 yields, for any $\delta\in(0,1]$,
@@ 6.4.4 Good-set STE inequality and explicit loop-length truncation
Fix truncation thresholds $\alpha_{i}\in(0,\alpha_{0}]$. Let $G(h,T;\alpha)$ be the event that all spikes up to $T$ satisfy $A_{i}^{k}\geq\alpha_{i}$ and that spike counts match at the observation time (no horizon mismatch), i.e. $N_{i}^{h}(T)=N_{i}(T)$ for every neuron ii.
@@
Under Assumption 6.11, the remainder term $G^{K+1}\Delta$ is absent on the good set once $K\geq K_{\mathrm{eff}}(T)$: each multiplication by $G$ corresponds to taking one synaptic predecessor, so $(G^{K+1}\Delta)_{i}$ expands into contributions along directed, time-ordered influence chains of length $K+1$ ending at neuron ii. By definition of $K_{\mathrm{eff}}(T)$, such chains do not exist on $G(h,T;\alpha)$ for $K=K_{\mathrm{eff}}(T)$. Therefore, on $G$ the Neumann series truncates and (104) holds. Taking $\|\cdot\|_{\infty}$ gives (105).
@@ 7.3 One-step weak defect: explicit dependence on $R_{\mathrm{net}}$ and $R_{\mathrm{net},2}$
7.4 Global weak order 11 with explicit rate dependence
Theorem 7.2 (Recurrent global weak order 11 with explicit $R_{\mathrm{net}},R_{\mathrm{net},2}$ scaling).
@@ 7.4 Global weak order 11 with explicit rate dependence
The bursty synchronous regime corresponds to large $R_{\mathrm{net}}(T)$ and $R_{\mathrm{net},2}(T)$. The weak order remains 11 under these bounds, but the constants can become large. Moreover, synchronized regimes can increase the strip-density constants $\rho^{V}_{\max,i}(T)$ and the velocity moments $A_{2,i}(T)$ (in small-noise limits these may blow up), further increasing $C_{\mathrm{weak}}^{\mathrm{rec}}(T,W)$.